# One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings

### v1

@@ Appendix A Synthetic dataset
SynthCode2Code2NL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
@@
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCode2Code2NL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 5, In Figure 6 and Figure 7 some examples of code translation are shown.
@@
Figure 7: Example of a code snippet translation from Python to PHP.

### v2

@@ Appendix B Synthetic dataset
SynthCoNL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
@@
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCoNL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 11, In Figure 12 and Figure 10 some examples of code translation are shown.
Table 4 shows the average number of characters per language in SynthCoNL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in the Limitations section.
Appendix C Failure Case Analysis in Code-to-Code Retrieval
Figure 7: Retrieved incorrect samples from CodeT5+, UniXcoder, and the 18th layer of MoSE. All the models appear to be misled by structural similarities, such as loop functions and return values, highlighting a common superficial retrieval phase.
Figure 8: Histograms of normalized similarity score between correct and incorrect retrieved code snippets. The first 5 distributions belong respectively to MoSE 4, 9, 18, 27, 36 layers’ exit points; CodeT5+ and UniXcoder are the last ones displayed. While our model tends to have different, distinct distributions regarding wrong and correct retrievals, both UniXcoder and CodeT5+ have more overlapping results.
Table 5: Permutation tests between correct and incorrect similarity scores ($\alpha=0.05$). The table shows a big discrepancy between MoSE t-statistics and both UniXcoder and CodeT5+ results.
Model t-statistic p-value
MoSE (L4) 8.0 <0.001
MoSE (L9) 7.3 <0.001
MoSE (L18) 8.0 <0.001
MoSE (L27) 7.8 <0.001
MoSE (L36) 6.3 <0.001
UniXcoder 2.2 0.04
CodeT5+ 110M-embedding 2.7 0.009
Figure 9: Heatmap of permutation tests (10,000 permutations) between MoSE similarity score results ($\alpha=0.05$). * with p<0.05 and ** with p<0.001. In the left heatmap, we present the differences in correctly retrieved snippets across various layers, while the right heatmap illustrates the incorrectly retrieved snippets. The heatmaps demonstrate that even layers trained to achieve similar representation objectives can yield different similarity scores.
In this section, we analyzed where our model fails in code-to-code retrieval, relying upon gfg, a Python-to-Java dataset. We compared how our model gets misled by different code snippets with CodeT5+ and UniXcoder, and the confidence level with which the model identifies the occurrences. In fig. 7 we highlight a common defect of the most used open code encoders. When retrieving code snippets focusing on semantic similarity, we emphasize how all the inspected models might get misled by structural factors such as loops, if statements, and return types. The retrieved results underline an overconsideration of structural features despite semantic differences.
We then analyzed whether the model’s confidence might suggest uncertainty in retrieving incorrect snippets of code, indicating a statistically significant level of "awareness" in overconsidering misleading features. Thus, we compared correct and incorrect similarity distribution scores for retrieved code snippets. Figure 8 shows how MoSE, CodeT5+, and UniXcoder results differ when comparing correct and incorrect retrievals. In table 5, MoSE demonstrated a more significant t-statistic than CodeT5+ and UniXcoder. In a retrieval use case scenario, when the confidence level is insufficient, we suggest that thresholding the similarity score could be an option when deciding whether to provide the end user with a single result or a herd of results.
In fig. 9, we present the differences in similarity scores among the five exit points using a permutation test. It is important to note that, although the model is trained to maintain similar retrieval capabilities across these five exit points (as discussed in section 2.1), and while the incorrectly retrieved snippets might overlap between the layers, each exit point encodes code snippets differently.
Based on these results, we strongly recommend conducting a grid search between the similarity score threshold and the exit point to utilize.
Figure 10: Example of a code snippet translation from Python to PHP.
@@ Appendix C Failure Case Analysis in Code-to-Code Retrieval
Appendix D Model card
Table 6 provides a comprehensive overview of the hyperparameters, architectural choices, and training configurations used in the development of MoSE. This includes detailed settings for both the pre-training and fine-tuning stages, as well as hardware specifications. The information is provided to enhance reproducibility and improve understanding of the model’s design.
Table 6: Details on hyperparameters, architecture, and training for MoSE are presented here, including insights from both pre-training and fine-tuning stages.
Architecture Pre-training
Hyperparameter Value Hyperparameter Value
Model size 1B parameters Batch size 3.99M tokens
Precision bfloat16 Pretraining steps 245,000
Hidden layers 36 Pretraining tokens 1T
Attention heads 16 Loss function MLM + In-Context loss
Hidden dimensionality 1,024 Multi-layer loss Yes
Positional encoding RoPE ($\theta=10^{-6}$) Optimizer AdamW
Context length 2,048 Weight decay $1\mathrm{e}{-1}$
Attention mechanism Grouped Query Attention Initial learning rate $6.24\mathrm{e}{-4}$
Attention pattern Bi-directional Learning rate schedule Multi-step
Warmup steps 4,000
Fine-tuning Hardware (Pre-training + Fine-tuning)
Hyperparameter Value Hyperparameter Value
Dataset size 635,404 samples Pre-training GPUs 512 NVIDIA Ampere A100 (64GB)
Fine-tuning steps 20,000 Fine-tuning GPUs 512 NVIDIA Ampere A100 (64GB)
Loss function CLIP loss Benchmarking GPUs 1 NVIDIA Ampere A100 (64GB)
Multi-layer loss Yes Overall training hours 450,000
Batch size 2,048
Learning rate $1.0\mathrm{e}{-5}$
Temperature parameter 10.0